OpenStudy (anonymous):

Simplifying Radicals ; Square Root Of W^8

jimthompson5910 (jim_thompson5910):

Hint: \[\Large \sqrt{x^k} = x^{\frac{k}{2}}\]

jimthompson5910 (jim_thompson5910):

Basically, you divide the exponent by 2

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

np